Part D Calculate the mass defect of the helium nucleus 32He. The mass of neutral 32He...

Part D

Calculate the mass defect of the helium nucleus 32He. The mass of neutral 32He is given by MHe=3.016029amu.

Express your answer in atomic mass units to four significant figures.

Part E

Calculate the binding energy E of the helium nucleus 32He (1eV=1.602�10?19J).

Express your answer in millions of electron volts to four significant figures.

E =   

Part F

Calculate the binding energy per nucleon of the helium nucleus 32He.

Express your answer in millions of electron volts to four significant figures.
Answer #1

D)

given 3 He 2

we know that

number of protons = atomic number = 2

also

number of neutrons = mass number - atomic number

so

number of neutrons = 3 -2 = 1

now

mass of nucleus = ( 2 x mass of proton ) + ( 1 x mass of neutron)

so

mass of nucleus = ( 2 x 1.007825u) + ( 1.008665u)

mass of nuclues = 3.024315u

now

mass defect (dm) = 3.024315 - 3.016029

mass defect (dm) = 8.286 x 10-3 u

so

the mass defect is 8.286 x 10-3 u


E)

now

we know that

binding energy = dm x 931.5 MeV / u

so

B.E = 8.286 x 10-3 u x 931.5 MeV / u

B.E = 7.718409 MeV

so

the binding energy is 7.718 MeV


F) now

number of nucleon = 3

so

binding energy per nucleon = 7.718 / 3

binding energy per nucleon = 2.573

so

the binding energy per nucleon is 2.573 MeV / nucleon